OpenStudy (anonymous):

Find the sum of the first 30 terms of the sequence below. 3n+2

OpenStudy (welshfella):

i'm assuming that 3n + 2 is the nth term so when n=1 the first term is 3(1)+2 = 5

OpenStudy (welshfella):

sorry that's 3(2) + 2 = 8

OpenStudy (anonymous):

thanks

OpenStudy (welshfella):

3rd term is 3(3) + 2 = 11 first 3 terms are 5,8,11 what type of sequence is this?